In 2018, Brady purchases a 2018 Nissan Leaf electric vehicle for his personal use. He is eligible to claim a credit of $7,500. He is in the 35 percent marginal tax bracket and his regular tax liability before credits is $14,800. What is the tax benefit Brady realizes from this purchase?
Answer #1

Hello there !!

To understand this easily, look at the base of the question, it says Brady is in the 35% marginal tax bracket and his regular tax liability before the eligible credits of $7500 is $14800. Because the income is not given in the question we need to make back calculation to compute the income first, then adjust the eligible credit and recompute the taxes and then we'll be able to get the tax benefit he realizes.

Calculation of Income before claiming credit on which tax was computed

$14800 / 35 * 100 = $42,286

Now claim the eligible credit from the income which will give you the income after claiming credits

$42,286 - $7500 = $34786

Now on the income after claiming credits compute the tax liability

$34786 * 35/100 = $12175

Therefore the tax benefit Brady realizes from this purchase is

$14800 - $12175 = $2625.

It can also be calculated in a shortcut method as tax saving on the eligible credit that on the claim of $7500 directly tax can be charged which is the net tax saving i.e., $7500*35/100 = $2625

Hope this clears your question. In case of any queries please feel free to ask.

And don't forget to leave your valuable feedback. :)